If a person received a test score that is in the top 32% of all test scores, the person’s z-score must be at least 0.46. How did we come up with 0.46? What is the formula used?
Answer #1

Solution:-

Given that

Using standard normal table,

P(Z > z) = 32%

= 1 - P(Z < z) = 0.32  

= P(Z < z) = 1 - 0.32

= P(Z < z ) = 0.68

= P(Z < 0.46 ) = 0.68  

z = 0.46